# Towards Interpretable and Efficient Attention: Compressing All by Contracting a Few

### v2

@@ Abstract
Attention mechanisms in Transformers have gained significant empirical success. Nonetheless, the optimization objectives underlying their forward pass are still unclear. Additionally, the quadratic complexity of self-attention is increasingly prohibitive. Unlike the prior work on addressing the interpretability or efficiency issue separately, we propose a unified optimization objective to alleviate both issues simultaneously. By unrolling the optimization over the objective, we derive an inherently interpretable and efficient attention mechanism, which compresses all tokens into low-dimensional structures by contracting a few representative tokens and then broadcasting the contractions back. This Contract-and-Broadcast Self-Attention (CBSA) mechanism can not only scale linearly but also generalize existing attention mechanisms as its special cases. Experiments further demonstrate comparable performance and even superior advantages of CBSA on several visual tasks. Code is available at this https URL.
@@ 1 Introduction
The Transformer architecture and attention mechanisms [1] have been widely applied across diverse areas, including computer vision [2, 3], natural language processing [4, 5], and scientific discovery [6]. Nonetheless, a series of puzzling phenomena have been uncovered in Transformers, such as the emergent segmentation properties [7], the in-context learning ability [8], the attention collapse [9, 10] and the extreme-token phenomena [11]. At the same time, the quadratic computational and memory complexity of self-attention with respect to the sequence length impedes its broader applications in real-time systems [12], as well as the processing of long documents [13] and high-resolution images [14].
@@
In light of these challenges, it is crucial to demystify attention mechanisms and enhance the efficiency of self-attention. Over the past few years, remarkable advances have been made in addressing the interpretability or efficiency issue separately. On one hand, attention mechanisms can be interpreted as unrolling an optimization objective grounded in clustering [15], compression [16], or energy minimization [17]. These ante-hoc, inherently interpretable perspectives are generally considered more rigorous than post-hoc explanations [18]. On the other hand, numerous techniques have been proposed to alleviate the quadratic complexity of self-attention, including sparse attention [19], and linear attention [20].
However, a unified principle that jointly achieves interpretability and efficiency in self-attention remains largely unexplored. This leaves the design of efficient attention mostly heuristic and the theories behind interpretable attention less instructive. To bridge this gap, a recent work [21] introduced the the maximal coding rate reduction (MCR2[22] objective but left several important questions unresolved. Their token-statistic self-attention (TSSA) adaptively scales feature channels based on their second moments, achieving high interpretability with linear computation complexity. Nevertheless, the channel attention [23, 24] manner of TSSA contrasts sharply with the spatial manner of standard softmax attention [1]. Therefore, it remains an open question whether an interpretable framework can be developed to integrate attention mechanisms, ranging from full attention  [^1] Softmax attention calculates all possible pairwise similarities among input tokens, i.e., it models a fully-connected digraph, and and is therefore also called full attention. to linear attention and channel attention, that trade expression capacity for increased efficiency while revealing their fundamental differences.
In this paper, we propose a unified optimization objective to derive an inherently interpretable and efficient attention mechanism. To be specific, the proposed objective compresses input tokens into a union of subspaces by contracting a few representative tokens (abbreviated as representatives), making it more efficient than processing all tokens individually. By unrolling the optimization procedure over the objective, we derive our Contract-and-Broadcast Self-Attention (CBSA), which contracts a few representatives and broadcasts the contractions back to all input tokens, as illustrated in Fig. 1.
Given a fixed number of representatives, CBSA’s computational and memory complexity scales linearly with the number of tokens. Moreover, varying the choice of representatives yields CBSA variants that trade efficiency for expression capacity, or vice versa. We demonstrate that: a) when each token serves as its own representative, CBSA recovers softmax attention; b) when representatives are orthogonal (e.g., the principal directions), CBSA reduces to linear attention; c) when representatives are orthogonal but fixed (i.e., input-agnostic), CBSA degenerates into channel attention. Our findings suggest that differences among these attention mechanisms arise from distinct choices of representatives, whose number and structure induce divergent patterns of information propagation.
Figure 1: Contract-and-Broadcast Self-Attention (CBSA). Besides projection onto the subspace and back to the ambient space, our CBSA generally involves four steps: initialization, extraction, contraction, and broadcast. In Section 3.3, we show that different representative extraction strategies can simplify or fuse some of these steps, thereby leading to different attention mechanisms. A more easy to understand
@@
    We propose a unified optimization objective that compresses all tokens by contracting a few representatives, thereby deriving an inherently interpretable and efficient attention mechanism.
@@
    By varying the choice of representatives, the derived mechanism, CBSA, generalizes existing attention mechanisms as special cases, reducing their fundamental differences to the number and structure of representatives.
@@
    We validate the interpretability and efficiency of CBSA through extensive experiments on visual tasks, highlighting its competitive or superior performance.
@@ 2 Notations and preliminaries
Given a positive integer $n$, we denote $n]\doteq\{1,2,\ldots,n\}$. For $s\geq t$, we use $\mathcal{O}(s,t)\subseteq\mathbb{R}^{s\times t}$ to denote the set of $s\times t$ matrices with orthonormal columns, and $\mathcal{O}(s)\doteq\mathcal{O}(s,s)$ to denote the set of $s\times s$ orthogonal matrices. We denote an identity matrix of size $n$ by $\mathbf{I}_{n}$, and similarly a zero square matrix by $\mathbf{O}_{n}$. Given a vector $\boldsymbol{v}\in\mathbb{R}^{n}$, let $\operatorname{Diag}(\boldsymbol{v})\in\mathbb{R}^{n\times n}$ be a diagonal matrix with the elements of $\boldsymbol{v}$ along the diagonal.
We use $\boldsymbol{Z}\in\mathbb{R}^{d\times N}$ to denote a sequence of $N$ input tokens represented in the $d$-dimensional ambient space, and specify these fed into the $\ell$-th attention layer as $\boldsymbol{Z}^{\ell}$. The same holds for the tokens’ representatives $\boldsymbol{Q}\in\mathbb{R}^{d\times m}$, where $m$ can be much smaller than $N$. The low-dimensional structure of representation is modeled as $K$ incoherent $p$-dimensional subspaces spanned by orthonormal bases $\boldsymbol{U}_{[K]}\doteq\{\boldsymbol{U}_{k}\in\mathcal{O}(d,p)\}_{k=1}^{K}$, such that $pK=d$ and $\boldsymbol{U}_{i}^{\top}\boldsymbol{U}_{j}=\mathbf{O}_{p},\forall i\neq j$. For the $i$-th basis vector of $\boldsymbol{U}_{k}$, we denote by $\boldsymbol{u}_{ki}$. Similar to [16], we implement $\boldsymbol{U}_{[K]}$ as learnable parameters in each attention layer, denoting the set modeled by the $\ell$-th layer as $\boldsymbol{U}^{\ell}_{[K]}=\{\boldsymbol{U}_{k}^{\ell}\}_{k=1}^{K}$.
To quantify compactness, i.e., the extent to which the tokens are compressed towards the subspaces, we adopt the (lossy) coding rate [25], which measures how efficiently the token distribution can be covered by $\epsilon$-balls under a given quantization precision $\epsilon>0$. The coding rate of the input tokens in the ambient space is defined as:
$R(\boldsymbol{Z})\doteq\frac{1}{2}\log\det\left(\mathbf{I}_{N}+\frac{d}{N\epsilon^{2}}\boldsymbol{Z}^{\top}\boldsymbol{Z}\right),$ (1)
and similarly defined for the projections onto the $k$-th subspace as $R(\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})$.
@@ 3.1 Compressing all by contracting a few
Before presenting our proposed optimization objective, we start from a well-established learning objective grounded in compression, maximal coding rate reduction (MCR2[22, 26], and streamline it by introducing representatives.
The MCR2 objective aims to compress tokens into $K$ subspaces while expanding them in the ambient space to learn compact and structured representation. Specifically, it quantifies both compression and expansion through the coding rate, and can be formulated as follows:
$\max_{\boldsymbol{Z}}\ \Delta R(\boldsymbol{Z})\doteq\underbrace{R(\boldsymbol{Z}))}_{\text{expansion}}-\underbrace{R_{c}(\boldsymbol{Z}\mid\boldsymbol{U}_{[K]})}_{\text{compression}}\doteq R(\boldsymbol{Z})-\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}).$ (2)
By unrolling the optimization over the compression term of (2), Yu et al. [16] derived an inherently interpretable softmax attention mechanism called multi-head subspace self-attention (MSSA) (9), which also scales quadratically. This complexity essentially arises from the fact that (1) calculates the Gram matrix of tokens in order to compress them. [^2] By the commutative property of the coding rate [25], the Gram matrix can be equivalently replaced by the covariance matrix, which directly addresses the quadratic complexity [27, 21]. However, as discussed in the introduction, this leads to an invertible degeneration to channel attention.
Inspired by the scaling strategies built upon landmarks [28, 29, 30], in this paper, We propose to streamline the MCR2 objective by compressing all input tokens through contracting a small number of representatives of them, resulting in a new objective that requires minor modifications to the compression term of (2):
$\max_{\boldsymbol{Z}}\ R(\boldsymbol{Z})-\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})\quad\text{s.t.}\ \sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})=\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}),$ (3)
@@
where the optimization variable involves only input tokens $\boldsymbol{Z}$, as representatives $\boldsymbol{Q}$ is extracted from it. [^3] Note that the representatives are not necessarily a subset of the input tokens; rather, they resemble the cluster centroids in $K$-means and are continuously determined, rather than discretely selected from the inputs. When the constraint of (3) holds, (3) is equivalent to (2) but with much higher efficiency, because the number of representatives (e.g., $m=p=\nicefrac{{d}}{{K}}$) is usually far smaller.
Relaxation and regularization of (3). As the equality constraint of (3) is overly restrictive in practice, we relax it by introducing a tolerance parameter $\tau$, which bounds the absolute difference between the two coding rates within each subspace, resulting in the following inequality constraints:
$\left|R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})\right|\leq\tau,\ \forall k\in[K].$ (4)
As a result, contracting the representatives via (3) ensures that the input tokens are correspondingly compressed in most cases, despite a potential significant gap between their coding rates. [^4] In Fig. 6, we observe that although the gap (i.e., $\tau$) is significant, the trends of the two coding rates still closely echo each other. We thus do not deliberately constrain the value of $\tau$. Additionally, note that the coding rate is invariant under any orthogonal transformation [22]. In extreme cases, if no additional structural constraint linking representatives to input tokens is imposed, the representatives could even be orthogonal to all input tokens, while still yielding identical coding rates. Therefore, we regularize the relaxed constraints (4) through a reconstruction term, such that the representatives can be expressed as linear combinations of the input tokens.
Eventually, the optimization objective that our subsequent derivations rely on is formulated as:
$\min_{\boldsymbol{Z}}$ $\ \sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})$
s.t. $\ \left|R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})\right|\leq\tau,\ \boldsymbol{U}^{\top}_{k}\boldsymbol{Q}=\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}\boldsymbol{A}_{k},\ \forall k\in[K],$ (5)
where $\{\boldsymbol{A}_{k}\in\mathbb{R}^{N\times m}\}^{K}_{k=1}$ are the coefficient matrices over the dictionaries $\{\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\in\mathbb{R}^{p\times N}\}_{k=1}^{K}$. Following [16], we optimize the expansion term $R(\boldsymbol{Z})$ in alternation with the compression term via a feed-forward module named ISTA, and thus have omitted it in (5).
@@ 3.2 Contract-and-Broadcast Self-Attention
In this subsection, we derive our inherently interpretable and efficient attention mechanism by translating the gradient-based optimization procedure over (5) into forward-pass operations.
Representative initialization and extraction. According to [31], an inherently interpretable cross-attention mechanism based on MCR2, called multi-head subspace cross-attention (MSCA), can be interpreted as aligning the coding rates of its query and key/value  [^5] In the context of inherently interpretable attention mechanisms based on MCR2, the query, key, and value coincide in self-attention, while the key and value coincide in cross-attention. The reason for this will become clear in the following derivation of CBSA. Therefore, we begin with an initial guess of the representatives as the query of MSCA, and take the input tokens as the key and value to extract representatives that satisfy the constraints in (5). In addition, the attention matrices in the attention heads of MSCA  [^6] That is, $\boldsymbol{A}_{k}=\operatorname{softmax}\left((\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})^{\top}(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q}_{0})\right)$, where $\boldsymbol{Q}_{0}$ denotes the initialized representatives. directly serve as the coefficient matrices in (5), since the query is updated as linear combinations of the key. Following [30], we employ a pooling-based strategy for representative initialization. [^7] Another intuitive choice is to initialize them as learnable parameters, as in [32]. However, we find that this strategy suffers from a low-rank problem, which is discussed in the Appendix A. More details are deferred to the Appendix A, as our primary focus here is on the subsequent derivation.
Representative contraction and contraction broadcast. Having initialized the representatives via pooling the input tokens and extracted those satisfying the constraints of (5) via MSCA, we move on to taking a gradient descent step on the objective of (5):
$\boldsymbol{Z}\leftarrow\boldsymbol{Z}-\eta\nabla_{\boldsymbol{Z}}\left(\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})\right)=\boldsymbol{Z}-\kappa\operatorname{CBSA}(\boldsymbol{Z}\mid\boldsymbol{U}_{[K]}),$ (6)
@@
where $\eta$ denotes the original step size of the gradient descent update, and the step-size $\kappa$ is introduced to absorb the gradient coefficient for notational simplicity. The contraction term in (7) is named so because it is proportional to the gradient of the representatives’ coding rate with respect to themselves, which specifies the contraction direction (abbreviated as contractions) of them. The broadcast term in (7), an $m\times N$ matrix encoding the linear relationships between representatives and input tokens, serves to broadcast the contractions back to all input tokens.
To reduce the computational overhead of the contraction step, we adopt [16, Approximation 4], which replaces the matrix inverse with a Gram matrix followed by a softmax (i.e., an attention matrix), yielding the CBSA operator implemented in our Transformer models: [^8] Note that the scaling factor and possible sign inversion introduced by this approximation are absorbed by $\kappa$ again, which is implemented as a learnable parameter in the model.
@@
The self-attention operation above corresponds to softmax attention, except that the linear projections for query, key, and value are all replaced by the subspace basis, i.e., $\boldsymbol{W}^{\mathrm{q}}=\boldsymbol{W}^{k}=\boldsymbol{W}^{v}=\boldsymbol{U}_{k}^{\top}$. By default, we adopt this simplified version (8) in our Contract-and-Broadcast Transformer (CBT), as illustrated in Fig. 1. Nonetheless, as will be shown in the next subsection, it is the rigorous prototype (7) that exposes the fundamental connections among different attention mechanisms.
Figure 2: FLOPs of a single attention layer, with $d=384$, $H=6$, and a patch size of $16\times 16$.
Overview and complexity analysis of CBSA and CBT. Our CBSA efficiently processes input tokens through the following steps: it projects input tokens into subspaces (i.e., attention heads parameterized by $\boldsymbol{U}_{[K]}$), initializes representatives via pooling, extracts representatives via cross-attention, contracts them via self-attention, and broadcasts the contractions back to all tokens according to the attention matrices computed in the extraction step. For an overall interpretability, we adopt ISTA [16] as the feedforward block, as it optimizes the expansion term omitted from (5). We present the computational complexities of CBSA and its sub-operations in Table 1, and compare the FLOPs of different attention mechanisms in Fig. 2. Provided $N>\nicefrac{{2d}}{{H}}=2p$ (typically 128 in Transformers), the FLOPs of CBSA are lower than those of MSSA. Comparisons with more modules, such as MHSA and MLP, can be found in Fig. 9.
Table 1: Computational complexities. By default, we set $m=p=\nicefrac{{d}}{{H}}$, where $H=K$ denotes the number of attention heads (interpreted as subspaces in our case). The complexity of each sub-operation is computed by summing the costs across all heads. It is worth noting that the projection operations, which are essential to almost all attention mechanisms, bottlenecks the overall complexity always over $\mathcal{O}(Nd^{2})$.

### v3

@@ Abstract
Attention mechanisms have achieved significant empirical success in multiple fields, but their underlying optimization objectives remain unclear yet. Moreover, the quadratic complexity of self-attention has become increasingly prohibitive. Although interpretability and efficiency are two mutually reinforcing pursuits, prior work typically investigates them separately. In this paper, we propose a unified optimization objective that derives inherently interpretable and efficient attention mechanisms through algorithm unrolling. Precisely, we construct a gradient step of the proposed objective with a set of forward-pass operations of our Contract-and-Broadcast Self-Attention (CBSA), which compresses input tokens towards low-dimensional structures by contracting a few representatives of them. This novel mechanism can not only scale linearly by fixing the number of representatives, but also covers the instantiations of varied attention mechanisms when using different sets of representatives. We conduct extensive experiments to demonstrate comparable performance and superior advantages over black-box attention mechanisms on visual tasks. Our work sheds light on the integration of interpretability and efficiency, as well as the unified formula of attention mechanisms. Code is available at this https URL.
@@ 1 Introduction
Attention mechanisms have been widely applied across diverse areas, including computer vision [1, 2], natural language processing [3, 4], and scientific discovery [5]. Nonetheless, a series of puzzling phenomena—such as emergent segmentation properties [6], in-context learning ability [7], attention collapse [8, 9] and extreme-token phenomena [10]—have been uncovered in them, hindering the principled and trustworthy development. At the same time, the quadratic computational and memory complexity of self-attention with respect to the sequence length impedes its broader applications in real-time systems [11], as well as the processing of long documents [12] and high-resolution images [13].
@@
In light of these challenges, it has been more crucial to mathematically demystify attention mechanisms, which offers deeper insights into their simplification and acceleration. Over the past few years, remarkable advances have been made in addressing the interpretability or efficiency issue separately. On the one hand, in an ante-hoc manner, attention mechanisms can be interpreted by optimization objectives grounded in clustering [14], denoising [15], energy minimization [16], matrix decomposition [17], and contrastive learning [18]. These inherently interpretable approaches are more rigorous than post-hoc explanations [19]. On the other hand, numerous techniques have been developed to alleviate the quadratic complexity of self-attention, including sparse attention [20] and linear attention [21].
However, the joint development of interpretability and efficiency in attention mechanisms remains a largely unexplored area of research. This leaves the design of efficient attention mostly heuristic, and the interpretations and explanations for attention mechanisms less instructive. To bridge this gap, we formulate a unified optimization objective by mildly modifying a compression-driven optimization objective called MCR2 [22]. Indeed, this objective has been utilized for designing an interpretable softmax attention, MSSA [23], and a linear-time attention, TSSA [24]. But MSSA also scales quadratically, and TSSA is effectively a channel attention mechanism [25, 26], which contrasts sharply with both softmax attention (token mixer) and linear attention (channel mixer). [^1] Softmax attention calculates all possible pairwise similarities between tokens, while TSSA just scales feature channels according to their second moments. Therefore, instead of an isolated mechanism, we aim to develop a framework that unifies these varied attention mechanisms in an interpretable way, revealing how they are fundamentally connected yet distinctly presented, as well as the trade-off between expressive capacity and efficiency.
In this paper, we adopt two ante-hoc interpretations to constitute our proposed optimization objective: a) input tokens are compressed towards low-dimensional structures for compact and structured representation; and b) the geometry and information-theoretic essence of input tokens can be captured by a small number of representatives [27, 28] of them. Since the former has been formulated as the MCR2 objective [22, 23] (see Section 2), the remaining task is to leverage the representatives to optimize it, thereby efficiently compressing all by contracting a few (see Section 3.1). By unrolling the resulting optimization objective, we derive our Contract-and-Broadcast Self-Attention (CBSA), which contracts the representatives and broadcasts the contractions back to input tokens (see Section 3.2).
Given a fixed number of representatives, the computational and memory complexity of CBSA scales linearly with the number of input tokens. Moreover, CBSA covers the instantiations of varied attention mechanisms, including softmax attention, linear attention, and channel attention, when using different sets of representatives (see Section 3.3). As a result, CBSA serves as a unified formula for these attention mechanisms, and attributes their differences to their distinct information propagation (more precisely, compression) patterns induced by the different number and structure of representatives.
Figure 1: Overview of CBSA. Left panel: A detailed illustration of the operations in the proposed CBSA. Besides projecting tokens onto subspaces and back the ambient space, there are generally two stages in CBSA: 1) representative initialization and extraction; 2) representative contraction and contraction broadcast. The former extracts representatives satisfying the inequality constraints in (4), while the latter is a gradient step of the compression term in (4). Right panel: CBSA covers instantiations of varied attention mechanisms. Their compression patterns are distinct as illustrated above. Further analysis is elaborated in Section $3.$
@@
    We formulate an optimization objective that unifies the interpretability and efficiency of attention mechanisms through the idea of compressing all by contracting a few.
@@
    We derive an inherently interpretable and efficient attention mechanism, CBSA, which is a potential unified formula for different attention mechanisms.
@@
    We validate the interpretability and efficiency of CBSA through extensive experiments on visual tasks.
@@ 2 Notations and preliminaries
Notations. Given a positive integer $n$, let $n]\doteq\{1,2,\ldots,n\}$. For $s\geq n$, let $\mathcal{O}(s,n)\subseteq\mathbb{R}^{s\times n}$ denote the set of $s\times n$ matrices with orthonormal columns, and $\mathcal{O}(s)\doteq\mathcal{O}(s,s)$ denote the set of $s\times s$ orthogonal matrices. Let $\mathbf{I}_{n}$ denote an identity matrix of size $n$, and $\mathbf{O}_{n}$ denote a zero square matrix of size $n$. Given a vector $\boldsymbol{v}\in\mathbb{R}^{n}$, let $\operatorname{Diag}(\boldsymbol{v})\in\mathbb{R}^{n\times n}$ be a diagonal matrix with the entries of $\boldsymbol{v}$ along its diagonal. Let $\boldsymbol{Z}\in\mathbb{R}^{d\times N}$ denote $N$ input tokens represented in the ambient space $\mathbb{R}^{d}$. Specially, let $\boldsymbol{Z}^{\ell}$ denote these tokens feeding into the $\ell$-th attention layer. The representative tokens $\boldsymbol{Q}\in\mathbb{R}^{d\times m}$ (hereafter referred to as representatives) are extracted from the input tokens $\boldsymbol{Z}$ through a differentiable mapping $q(\cdot):\mathbb{R}^{d\times N}\to\mathbb{R}^{d\times m}$, where $m$ can be much smaller than $N$.
Union of subspaces. Although the union of nonlinear manifolds provides a better approximation [29], we adopt a much simpler structure: the union of (low-dimensional) linear subspaces. [^2] As this structure faces challenges in adapting to all modalities and tasks, we excluded natural language processing tasks from our experiments. But we argue that it serves as a feasible starting point for finer-grained structures; further discussion is available on our OpenReview page. Specifically, it is parameterized as $K$ incoherent $p$-dimensional subspaces spanned by orthonormal bases $\boldsymbol{U}_{[K]}\doteq\{\boldsymbol{U}_{k}\in\mathcal{O}(d,p)\}_{k=1}^{K}$, where $pK=d$ and $\boldsymbol{U}_{i}^{\top}\boldsymbol{U}_{j}=\mathbf{O}_{p},\forall\,i\neq j$. We refer to the $i$-th basis vector of $\boldsymbol{U}_{k}$ as $\boldsymbol{u}_{ki}$. Similar to [23], we implement $\boldsymbol{U}_{[K]}$ as learnable parameters in each attention layer and thus denote the $\boldsymbol{U}_{[K]}$ implemented by the $\ell$-th layer as $\boldsymbol{U}^{\ell}_{[K]}=\{\boldsymbol{U}_{k}^{\ell}\}_{k=1}^{K}$.
Coding rate. To quantify the compactness, i.e., the extent to which tokens are compressed towards subspaces, we adopt the (lossy) coding rate [30], which measures how efficiently the token distribution can be covered by $\epsilon$-balls under a given quantization precision $\epsilon>0$ (as illustrated in Fig. 1(a)). The coding rate of input tokens in the ambient space $\mathbb{R}^{d}$ is defined as:
$R(\boldsymbol{Z})\doteq\frac{1}{2}\log\det\left(\mathbf{I}_{N}+\frac{d}{N\epsilon^{2}}\boldsymbol{Z}^{\top}\boldsymbol{Z}\right).$ (1)
MCR2 objective. The Maximal Coding Rate Reduction (MCR2[22] objective adopted in [23] is defined on the coding rate as follows:
$\max_{\boldsymbol{Z}}\ \Delta R(\boldsymbol{Z}\mid\boldsymbol{U}_{[K]})\doteq\underbrace{R(\boldsymbol{Z})}_{\text{expansion}}-\underbrace{R_{c}(\boldsymbol{Z}\mid\boldsymbol{U}_{[K]})}_{\text{compression}}-\underbrace{\lambda\|\boldsymbol{Z}\|_{0}}_{\text{sparsity}}\doteq R(\boldsymbol{Z})-\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})-\lambda\|\boldsymbol{Z}\|_{0}.$ (2)
The input tokens are compressed towards $K$ subspaces by the compression term, while expanded in the ambient space by the expansion term to avoid collapse, yielding compact and structured representation [22]. Yu et al. [23] have demonstrated that the approximated gradient step of the compression term in (2) corresponds an interpretable softmax attention mechanism.
@@ 3.1 Compressing all by contracting a few
Due to the existence of Gram matrix in (1), the attention mechanism derived from (2), which is called Multi-head Subspace Self-Attention (MSSA), inevitably scales quadratically with the number of input tokens. While previous work has bypassed this issue by replacing the Gram matrix with the covariance matrix [31] and further introduced a variational formulation [24], these strategies degenerate the token mixer into a channel mixer and channel attention, respectively.
In this paper, inspired by the concept of landmarks [32, 33], we propose a simple but flexible approach to streamline the optimization of MCR2: compressing all input tokens by contracting a small number of representatives of them. Before demonstrating that this achieves linear complexity in $N$ and prevents the aforementioned degeneration, we first formulate it as a new optimization objective.
An initial attempt is to impose a set of equality constraints on the coding rates as follows:
$\max_{\boldsymbol{Z}}\ R(\boldsymbol{Z})-\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-\lambda\|\boldsymbol{Z}\|_{0}\ \ \text{s.t.}\ \ R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})=R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z}),\ \forall\,k\in[K],$ (3)
@@
where $\boldsymbol{Q}\doteq q(\boldsymbol{Z})$, and the exact formulation of $q(\cdot)$ will be introduced shortly.  [^3] The representatives are not necessarily a subset of the input tokens; they resemble the cluster centroids in $k$-means, which are continuously extracted, rather than discretely selected. This new objective in (3) is equivalent to the original objective in (2) but is more efficient to handle because the number of representatives (e.g., $m=p=\nicefrac{{d}}{{K}}$) can be far smaller.
Since that the equality constraints in (3) is unnecessarily restrictive in practice, we attempt to relax them by introducing a tolerance $\tau$, which uniformly bounds the absolute difference of the two coding rates within each subspace, i.e., $|R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})|\leq\tau$. Therefore, contracting the representatives will correspondingly compress the input tokens as well up to the tolerance $\tau$.
Consequently, we have a relaxed optimization problem for our subsequent derivations as follows:
$\max_{\boldsymbol{Z}}\ R(\boldsymbol{Z})-\sum_{k=1}^{K}R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-\lambda\|\boldsymbol{Z}\|_{0}\ \ \text{s.t.}\ \ |R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Q})-R(\boldsymbol{U}^{\top}_{k}\boldsymbol{Z})|\leq\tau,\ \forall\,k\in[K].$ (4)
In this paper, we employ an arguably simplest way to extract $\boldsymbol{Q}$ in each subspace: $\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\boldsymbol{A}_{k}$, where $\boldsymbol{A}_{k}\in\mathbb{R}^{N\times m}$ is the coefficient matrix over dictionary $\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\in\mathbb{R}^{p\times N}$, i.e., the projected representatives in each subspace are linear combinations of the projected input tokens.
@@ 3.2 Contract-and-Broadcast Self-Attention
We are now ready to derive the proposed attention mechanism in an ante-hoc interpretable manner by implementing a gradient step of the compression term in our objective as the forward pass. This methodology dates back its origins to the pioneering work [34], and is referred to as algorithm unrolling or unfolding [35].
Representative initialization and extraction. Inspired by the fact that cross-attention can be interpreted as approximating the coding rate [36], we employ this idea to extract representatives that satisfy the inequality constraints in (4), thus capturing the information-theoretic essence of input tokens. Specifically, we take an initial guess of the representatives as the query, and the input tokens as the key and value matrices, i.e.,
$\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}=\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}_{\text{ini}}+\eta\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}\underbrace{\operatorname{softmax}\left((\boldsymbol{U}_{k}^{\top}\boldsymbol{Z})^{\top}(\boldsymbol{U}_{k}^{\top}\boldsymbol{Q}_{\text{ini}})\right)}_{\boldsymbol{A}_{k}},\ \forall\,k\in[K],$ (5)
where the initial guess $\boldsymbol{Q}_{\text{ini}}\doteq\operatorname{sg}(\operatorname{avgpool}(\boldsymbol{U}_{k}^{\top}\boldsymbol{Z}))$ is obtained via average pooling over the input tokens, the step size $\eta$ is learnable in our implementation, and $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator. The operator ensures that the initialization strategy does not affect the subsequent derivation, so the attention matrix in (5) effectively corresponds to the coefficient matrix $\boldsymbol{A}_{k}$.
Representative contraction and contraction broadcast. To derive the attention mechanism, following [23], we focus on optimizing the compression term via a gradient descent step. Given that the extracted representatives satisfy the inequality constraints, we perform a gradient descent step on the compression term of the objective in (4) with respect to the input tokens as follows:
$\boldsymbol{Z}\leftarrow\boldsymbol{Z}-\kappa\nabla_{\boldsymbol{Z}}R_{c}(\boldsymbol{Q}\mid\boldsymbol{U}_{[K]})=\boldsymbol{Z}-\kappa\frac{p}{m\epsilon^{2}}\operatorname{CBSA}(\boldsymbol{Z}\mid\boldsymbol{U}_{[K]})\ \ \text{where}\ \$(6)
@@
in which the step size parameter $\kappa$ is also learnable in our implementation. One can verify that the contraction term in (7) is proportional to the gradient of the compression term in (4) with respect to the representatives. Hence, we refer to this formula as a Contract-and-Broadcast Self-Attention (CBSA), reflecting that: a) the contraction term gives the contracting directions of the representatives (hereafter referred to as contractions); and b) the broadcast term, which reuses the attention matrix in (5), broadcasts the contractions back to all input tokens.
Contraction via self-attention. To avoid computing the expensive matrix inverse in (7), similar to [23], we approximate it by a Gram matrix and a softmax function (i.e., an attention matrix): [^4] The scaling factor and sign inversion [37] introduced by this approximation are absorbed into the learnable step size $\kappa$ for notational simplicity.
@@
Note that the contraction term now effectively constitutes a self-attention operation in which the linear projections for the query, key, and value matrices are all identical to the subspace basis, i.e., $\boldsymbol{W}_{\!\!\text{query}}=\boldsymbol{W}_{\!\!\text{key}}=\boldsymbol{W}_{\!\!\text{value}}=\boldsymbol{U}_{k}^{\top}$. By default, we implement CBSA via (8) rather than (7) in our experiments.
Figure 2: Computation complexity.
Overview of CBSA. The workflow of CBSA is illustrated in the left panel of Fig. 1. We also construct an inherently interpretable Contract-and-Broadcast Transformer (CBT) by stacking CBSA with the ISTA module [23], which is also derived via algorithm unrolling. [^5] Note that the ISTA module is designed to unroll the suitably-relaxed proximal gradient step to address the difference of the sparsity penalty and the expansion term, i.e., $\min\lambda\|\boldsymbol{Z}\|_{0}-R(\boldsymbol{Z})$. We report the computational complexities of CBSA and its sub-operations in Table 1 and compare the FLOPs of different attention mechanisms in Fig. 2 where $d=384$, $H=6$, and the patch size is $16\times 16$. Provided $N>\nicefrac{{2d}}{{H}}=2p$ (typically 128 in Transformers), the FLOPs of CBSA are lower than those of MSSA. Further comparisons with other modules, including MHSA [38] and MLP, are provided in Fig. 9.
Table 1: Computational complexities. By default, we set $m=p=\nicefrac{{d}}{{H}}$, where $H=K$ denotes the number of attention heads (interpreted as subspaces in our case). The complexity of each sub-operation is computed by summing the costs across all heads. It is worth noting that the projection operations, which are essential to almost all attention mechanisms, confine the overall complexity at least $\mathcal{O}(Nd^{2})$.